# Data-Driven Dynamics Modeling of Miniature Robotic Blimps Using Neural ODEs With Parameter Auto-Tuning

### v1

@@ Abstract
Miniature robotic blimps, as one type of lighter-than-air aerial vehicles, have attracted increasing attention in the science and engineering community for their enhanced safety, extended endurance, and quieter operation compared to quadrotors. Accurately modeling the dynamics of these robotic blimps poses a significant challenge due to the complex aerodynamics stemming from their large lifting bodies. Traditional first-principle models have difficulty obtaining accurate aerodynamic parameters and often overlook high-order nonlinearities, thus coming to its limit in modeling the motion dynamics of miniature robotic blimps. To tackle this challenge, this letter proposes the Auto-tuning Blimp-oriented Neural Ordinary Differential Equation method (ABNODE), a data-driven approach that integrates first-principle and neural network modeling. Spiraling motion experiments of robotic blimps are conducted, comparing the ABNODE with first-principle and other data-driven benchmark models, the results of which demonstrate the effectiveness of the proposed method.
@@
robotic blimp, dynamics modeling, data-driven method, first-principle method
@@ I Introduction
Aiming to precisely capture the complex dynamics of robotic blimps, this letter presents the Auto-tuning Blimp-oriented Neural Ordinary Differential Equation method (ABNODE), a neural ordinary differential equation model coupled with a data-driven algorithm that automatically adjusts both the physical and network parameters of the dynamics model. Specifically, this letter examines the spiral dynamics of robotic blimps, where the airflow field undergoes significant changes due to the blimp’s movement, resulting in uncertainties in aerodynamic parameters and non-negligible high-order nonlinear residual dynamics.
@@ II-A First-Principle Dynamics
where $A$ represents the reference area, determined by the blimp design. $K_{i}$ denotes the rotation damping coefficients. The terms in the form of $C_{\square}(\alpha,\beta)$ denote the aerodynamic coefficients typically modeled as polynomials of $\alpha$ and $\beta$ with the orders of no more than four [10].
Identifying the parameters in Eqs. (22) and (23) is crucial for accurately modeling robotic blimps. Traditionally, these physical parameters are derived through polynomial regression of experimental trajectory data obtained from selected motions such as straight-line flight or spiraling flight, with either constant or time-varying speeds. It is noteworthy that measuring these physical aerodynamic parameters via wind tunnel experiments is challenging and costly due to the small size and the flexible design of the envelope. Additionally, first-principle modeling tends to overlook high-order nonlinear dynamics, resulting in a parameterized model that lacks sufficient accuracy, especially during complex agile aerial maneuvers.
@@ II-B Data-Driven Modeling
$\dot{\bm{x}}=$ $\bm{f}(\bm{x},\bm{u},\bm{\eta})$ (24)
$\bm{\eta}=$ $[\bm{\eta}_{F_{i}},\bm{\eta}_{M_{i}},K_{i}]^{T}\;(i=1,2,3)$ (25)
where $\bm{x}$ and $\bm{u}$ represent the system states and the control inputs, respectively. $\bm{\eta}$ represents the overall physical parameters of the model. $\bm{\eta}_{F_{i}}$ and $\bm{\eta}_{M_{i}}$ represent the aerodynamic parameters $C_{F_{i}}$ and $C_{M_{i}}$ in Eq. (22) and Eq. (23), respectively. $K_{i}$ is the same damping parameter as in Eq. (23). The aim is to establish a continuous-time function, denoted as $\bm{f}$, to accurately model the dynamics given experimental flight data of the robotic blimp.
@@
Taking the advantages of both domain knowledge and data-driven modeling techniques, researchers have proposed various physics-informed neural networks for modeling the dynamics of complex systems. For example, Saviolo et al. introduced the physics-inspired temporal convolutional network (PI-TCN), which initially aligned with experimental data distribution and subsequently with first-principle model predictions in a two-phase process [15]. Jiahao et al. developed the knowledge-based neural ordinary differential equation (KNODE), integrating first-principle modeling with neural ordinary differential equations[19].
While PI-TCN and KNODE utilize both first-principle models and neural networks for dynamics modeling of robotic systems, they have limitations, especially concerning complex blimp dynamics. First, the model accuracy heavily depends on the initial setting of physical parameters. Inappropriate settings result in slow convergence or sometimes model failure. Second, these two models are typically trained to learn the system dynamics without considering the differences between kinematics and dynamics. In the case of robotic blimps, the first-principle kinematics model is often accurate, whereas the dynamics involve aerodynamic uncertainties, which both PI-TCN and KNODE overlook in their model training.
@@ III Design of the ABNODE Model
This section presents the architecture design of the ABNODE model, along with the physics-informed loss and the associated model training process.
@@ III-A First-Principle and Neural Network Modules
The ABNODE model consists of two modules including the first-principle module which incorporates the physical parameters of robotic blimps, and the neural network module, designed to capture residual dynamics. Given the accurate kinematics description (Eq. (7)) of the robotic blimp, the proposed ABNODE model only updates the dynamics (Eq. (17)), leaving the kinematics unchanged. Assuming the presence of additive high-order nonlinear uncertainty, we write the dynamics of the robotic blimp as
@@ III-C Two-phase Model Training
ABNODE features both trainable physical and network parameters along with a physics-informed loss function. This letter adopts a two-phase model training strategy, illustrated in Fig. 1. In the initial phase, $\bm{b}$ is set to 1 to prioritize updating the physical parameters $\bm{\eta}$, leveraging the potential of the first-principle module. Once convergence of the initial values of the physical parameters $\bm{\eta}^{\star}$ is achieved, $\bm{b}$ is then set to 0 to prioritize the updating of the neural network parameters $\bm{\theta}$. The proposed ABNODE is detailed in Algorithm 1. $\bm{X}=[\bm{x}_{0},\bm{x}_{1},\cdots,\bm{x}_{N}]$ represents the experimental data of system states from time step 00 to time step $N$. $\bm{\Upsilon}=[\bm{u}_{0},\bm{u}_{1},\cdots,\bm{u}_{N}]$ denotes the control inputs during the experiment. $\tilde{\bm{X}}=[\tilde{\bm{x}}_{0},\tilde{\bm{x}}_{1},\cdots,\tilde{\bm{x}}_{N}]$ and $\hat{\bm{X}}=[\hat{\bm{x}}_{0},\hat{\bm{x}}_{1},\cdots,\hat{\bm{x}}_{N}]$ represent the predicted states in the first phase and the second phase, respectively. The 4th-order Runge Kutta and Adam are selected as the ODE solver and the optimizer, respectively.
@@ III-C2 Second phase
3:  while not converged and $k<N_{1}+N_{2}$ do
4:     if $k>N_{1}$ then
5:        $\bm{b}\leftarrow 0$
6:     end if
7:     Solve Eq. (26) with the ODE solver
8:     Generate the prediction $\tilde{\bm{X}}$ or $\hat{\bm{X}}$ by Eqs. (29) or (30)
9:     Minimize the physical-informed loss (27), and calculate the gradient of $\bm{\eta}$ or $\bm{\theta}$
10:     $\bm{\eta}\leftarrow\bm{\eta}_{k+1}$ or $\bm{\theta}\leftarrow\bm{\theta}_{k+1}$
11:     $k\leftarrow k+1$
12:  end while
@@ IV-B Comparison Models
We conducted a comparative analysis of the ABNODE method against three benchmark models including the first-principle, KNODE, and SINDYc. The first-principle model followed Eqs. (7) and (17), with its aerodynamic parameters identified through solving the algebraic equations derived from (quasi-)steady flight experiments [10]. KNODE [23] incorporated the first-principle model into a neural network architecture. Unlike ABNODE, KNODE did not update the parameters of the first-principle model during training. SINDYc employed nonlinear candidate functions, including polynomials, sine and cosine functions, and their respective products (implemented via PySINDY [24]) to model the dynamics of the robotic blimp.
Figure 3: Heat maps of the prediction loss for the first-principle, SINDYc, KNODE, and ABNODE models, tested with spiral trajectories featuring various configurations. The vertical axis represents different displacements $\Delta_{\overline{\bm{r}}_{x}}$ of the gondola on the slide. The horizontal axis represents different configurations of the propulsion thrust $F_{l}$ and $F_{r}$.
@@ IV-C Model Configuration
Out of the 120 spiral experimental trajectories, we adopted a 3:1 train-test split. For every control input configuration, three spiraling trials were randomly selected for training the ABNODE and the two data-driven comparison models, while one spiraling trial was reserved for testing all four models. The performance of all models was evaluated using the loss function specified in Eq. (28). To normalize the data across various experimental configurations, we used the reciprocal of the range of the motion data as scaling factors, which served as the diagonal elements of the weight matrix $\bm{W}$.
@@ V-A Accuracy
We compared the prediction loss between ABNODE and other models. Figure 3 illustrates the heat maps of the averaged prediction loss across all the test trials concerning the control input configurations. We observe that ABNODE consistently outperforms the first-principle model, achieving a significant $63.58\%$ reduction in loss across all control input configurations. Furthermore, ABNODE demonstrates improvements of $53.81\%$ and $20.42\%$ in loss reduction compared to SINDYc and KNODE, respectively.
We adjusted the loss and designed two performance metrics, i.e., the time-dependent mean squared error (MSE) and time-dependent cumulative MSE (CMSE) of the prediction loss to delve deeper into the modeling accuracy over specific time periods. Let $\text{MSE}(t_{i})$ denote the model prediction error at time $t_{i}$ and $\text{CMSE}(t_{k})$ denote the sum of MSEs over $k$ time steps, i.e.,
@@
where $n$ represents the dimension of system states, and $\bm{x}(t_{i})$ and $\bm{x}_{\text{p}}(t_{i})$ represent the experimental data and model prediction of the states, respectively. As an illustration, Fig. 4 showcases the trajectory of CMSE and the distributions of MSE for the configuration $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(3\text{cm},1.7\text{gf},5.4\text{gf})$. ABNODE demonstrates the highest modeling accuracy throughout the trajectory compared to other models, evident from its smallest average MSE, indicating its temporal consistency and robustness.
Figure 5 shows the density histograms of the velocity and angular velocity, both from the experimental data and the model predictions by ABNODE. The similarity in distribution between the two sets validates the modeling accuracy concerning motion states from a statistical perspective.
Figure 5: Density histograms of velocity and angular velocity for both experimental and ABNODE-generated trajectories. Both velocity and angular velocity are expressed in the velocity frame $O_{v}-x_{v}y_{v}z_{v}$. The left column illustrates velocity distributions in $x_{v}$, $y_{v}$, and $z_{v}$ directions from top to bottom, while the right column shows angular velocity distributions in the same directions.
@@ V-B Generalizability
We evaluated the generalization capability of the four modeling methods using the spiraling experimental dataset. To do so, we trained models using one control input configuration and tested them on neighboring configurations For instance, the configuration $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(1\text{cm},1.2\text{gf},6.1\text{gf})$ had four neighboring configurations, as depicted in Fig. 3, including $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(1\text{cm},2.0\text{gf},5.0\text{gf})$, $(1\text{cm},1.7\text{gf},6.1\text{gf})$, $(0\text{cm},1.2\text{gf},6.1\text{gf})$, and $(2\text{cm},1.2\text{gf},6.1\text{gf})$. We excluded the edge configurations and selected 12 setups, each with four neighboring configurations, resulting in a total of 48 configurations for evaluating the model generalization capability.
Figure 6: Boxplot of the model prediction loss (Eq. 28) across four comparison models. The box represents the interquartile range (IQR), spanning from the first quartile (Q1) to the third quartile (Q3). The line inside the box represents the median while the square dot represents the mean. The whiskers extend from the box to the minimum and maximum values within 1.5 times the IQR.

### v2

@@ Abstract
Miniature robotic blimps, as one type of lighter-than-air aerial vehicles, have attracted increasing attention in the science and engineering community for their enhanced safety, extended endurance, and quieter operation compared to quadrotors. Accurately modeling the dynamics of these robotic blimps poses a significant challenge due to the complex aerodynamics stemming from their large lifting bodies. Traditional first-principle models have difficulty obtaining accurate aerodynamic parameters and often overlook high-order nonlinearities, thus coming to their limit in modeling the motion dynamics of miniature robotic blimps. To tackle this challenge, this letter proposes the Auto-tuning Blimp-oriented Neural Ordinary Differential Equation method (ABNODE), a data-driven approach that integrates first-principle and neural network modeling. Spiraling motion experiments of robotic blimps are conducted, comparing the ABNODE with first-principle and other data-driven benchmark models, the results of which demonstrate the effectiveness of the proposed method.
@@
Dynamics, calibration and identification, machine learning for robot control.
@@ I Introduction
Aiming to precisely capture the complex dynamics of robotic blimps, this letter presents the ABNODE, a neural ordinary differential equation model coupled with a data-driven algorithm that automatically adjusts both the physical and network parameters of the dynamics model. Auto-tuning refers to the automatic adjustment of aerodynamic parameters in the first-principle module of the proposed ABNODE. This letter examines the spiral dynamics of robotic blimps, where the airflow field undergoes significant changes due to the blimp’s movement, resulting in uncertainties in aerodynamic parameters and non-negligible high-order nonlinear residual dynamics.
@@ II-A First-Principle Dynamics
where $A$ represents the reference area, determined by the blimp design. $K_{i}$ denotes the rotation damping coefficients. The terms in the form of $C_{\square}(\alpha,\beta)$ denote the polynomials of $\alpha$ and $\beta$ with the orders of no more than four [10].
Identifying the parameters in Eqs. (22) and (23) is crucial for accurately modeling robotic blimps [16, 17]. Traditionally, these physical parameters are derived through polynomial regression of experimental trajectory data obtained from selected motions such as straight-line flight or spiraling flight, with either constant or time-varying speeds. It is noteworthy that measuring these physical aerodynamic parameters via wind tunnel experiments is challenging and costly due to the small size and the flexible design of the envelope. Additionally, first-principle modeling tends to overlook high-order nonlinear dynamics[18, 19], resulting in a parameterized model that lacks sufficient accuracy, especially during complex agile aerial maneuvers.
@@ II-B Data-Driven Modeling
$\dot{\bm{x}}=$ $\bm{f}(\bm{x},\bm{u};\bm{\eta})$ (24)
where $\bm{x}$ and $\bm{u}$ represent the system states and the control inputs, respectively. $\bm{\eta}=[\bm{\eta}_{F_{i}},\bm{\eta}_{M_{i}},\bm{\eta}_{K_{i}}]^{T}$ where $i=1,2,3$ represents the overall uncertain physical parameters of the model. $\bm{\eta}_{F_{i}}$ and $\bm{\eta}_{M_{i}}$ represent the aerodynamic parameters in polynominals $C_{F_{i}}$ and $C_{M_{i}}$ in Eq. (22) and Eq. (23), respectively. $\bm{\eta}_{K_{i}}$ represent the same damping parameters $K_{i}$ as in Eq. (23). The aim is to establish a continuous-time function, denoted as $\bm{f}$, to accurately model the dynamics given experimental flight data of the robotic blimp.
@@
Taking the advantages of both domain knowledge and data-driven modeling techniques, researchers have proposed various physics-informed neural networks for modeling the dynamics of complex systems. For example, Saviolo et al. introduced the physics-inspired temporal convolutional network (PI-TCN), which initially aligned with experimental data distribution and subsequently with first-principle model predictions in a two-phase process [15]. Jiahao et al. developed the knowledge-based neural ordinary differential equation (KNODE)[23], which integrates neural ordinary differential equations [24] with first-principle modeling.
While PI-TCN and KNODE utilize both first-principle models and neural networks for dynamics modeling of robotic systems, they have limitations, especially concerning complex blimp dynamics. First, the model accuracy heavily depends on the initial setting of physical parameters. Inappropriate settings result in slow convergence or sometimes model failure. Second, these models typically learn system dynamics without differentiating between robot kinematics and dynamics. For blimps, the kinematics model is often accurate, while dynamics involve aerodynamic uncertainties overlooked by PI-TCN and KNODE. ABNODE is particularly effective for: 1) establishing continuous-time ODE models; 2) incorporating dynamics with uncertain physical parameters; and 3) utilizing accurate robot kinematics.
@@ III-A First-Principle and Neural Network Modules
The ABNODE model consists of two modules including the first-principle module which incorporates the physical parameters of robotic blimps, and the neural network module, designed to capture residual dynamics. Given the accurate kinematics description (Eq. (7)) of the robotic blimp, the proposed ABNODE model only updates the dynamics (Eq. (17)), leaving the kinematics unchanged. Following the literature [25, 26, 27] and assuming an additive form of high-order nonlinear uncertainties, we write the dynamics of the robotic blimp as
@@ III-C Two-phase Model Training
ABNODE features both trainable physical and network parameters along with a physics-informed loss function. This letter adopts a two-phase model training strategy, illustrated in Fig. 1. In the initial phase, $\bm{b}$ is set to 1 to prioritize updating the physical parameters $\bm{\eta}$, leveraging the potential of the first-principle module. Once convergence of the initial values of the physical parameters $\bm{\eta}^{\star}$ is achieved, $\bm{b}$ is then set to 0 to prioritize the updating of the neural network parameters $\bm{\theta}$. The proposed ABNODE is detailed in Algorithm 1. $\bm{X}=[\bm{x}_{0},\bm{x}_{1},\cdots,\bm{x}_{N}]$ represents the experimental data of system states from time step 00 to time step $N$. $\bm{\Upsilon}=[\bm{u}_{0},\bm{u}_{1},\cdots,\bm{u}_{N}]$ denotes the control inputs during the experiment. $\tilde{\bm{X}}=[\tilde{\bm{x}}_{0},\tilde{\bm{x}}_{1},\cdots,\tilde{\bm{x}}_{N}]$ and $\hat{\bm{X}}=[\hat{\bm{x}}_{0},\hat{\bm{x}}_{1},\cdots,\hat{\bm{x}}_{N}]$ represent the predicted states in the first phase and the second phase, respectively.
@@ III-C2 Second phase
3:  while $k<N_{1}+N_{2}$ do
4:     Solve Eq. (25) with the ODE solver
5:     if $k<N_{1}$ then
6:        Generate the prediction $\tilde{\bm{X}}$ by Eq. (28)
7:     else
8:        Generate the prediction $\hat{\bm{X}}$ by Eq. (29)
9:     end if
10:     Minimize the physical-informed loss (26)
11:     if $k<N_{1}$ then
12:        calculate the gradient of $\bm{\eta}$, $\bm{\eta}\leftarrow\bm{\eta}_{k+1}$
13:     else
14:        calculate the gradient of $\bm{\theta}$, $\bm{\theta}\leftarrow\bm{\theta}_{k+1}$
15:     end if
16:     $k\leftarrow k+1$
17:  end while
@@ IV-A Experimental Setup
Figure 3: Heat maps of the prediction loss for the first-principle, SINDYc, NODE, KNODE, BNODE and ABNODE models, tested with spiral and linear trajectories featuring various motion configurations. The vertical axis represents different displacements $\Delta_{\overline{\bm{r}}_{x}}$ of the gondola on the slide. The left vertical axis corresponds to the left two columns, and the right vertical axis corresponds to the right five columns. The horizontal axis represents different configurations of the propulsion thrust $F_{l}$ and $F_{r}$.
@@ IV-B Comparison Models
We conducted a comparative analysis of the ABNODE method against five benchmark models including the first-principle, SINDYc, NODE, KNODE and BNODE. The first-principle model followed Eqs. (7) and (17), with its imprecise parameters, such as aerodynamic coefficients, identified through solving the algebraic equations derived from quasi-steady-state flight experiments [10] while precise parameters, such as mass and dimensions, are physically measured with high confidence levels. SINDYc employed nonlinear candidate functions, including polynomials, sine and cosine functions, and their respective products (implemented via PySINDY [31]) to model the residual dynamics of the robotic blimp. NODE uses the standard MLPs for modeling, relying solely on data-driven approaches without incorporating domain knowledge about system dynamics. KNODE [23] incorporates the first-principle model (encompassing both dynamics and kinematics) into the NODE architecture, while BNODE and ABNODE incorporate only the dynamics. Unlike ABNODE, both BNODE and KNODE do not update the parameters of the first-principle model during training. A comparison of these models is detailed in Table II.
TABLE II: Comparison between neural network-based models
NODE[24] KNODE[23]  BNODE ABNODE
Kinematics
Dynamics
Auto-tuning
@@ IV-C Model Configuration
Out of the 120 spiral and 20 linear experimental trajectories, we adopted a 3:1 train-test split. For every control input configuration, three trials were randomly selected for training the ABNODE and the four data-driven comparison models, while one spiraling trial was reserved for testing all six models. The network architecture of MLPs in the NODE model is consistent with that in the KNODE model for comparison purposes. The architecture includes a first layer with a dimension of $23\times 256$, a second layer of $256\times 64$, and a third layer of $64\times 12$. BNODE and ABNODE shared the same network structure as NODE and KNODE, except for their output layer, which had 6 dimensions corresponding to the dynamics output.
Regarding parameter initialization, the uncertainty parameters in the first-principles model, including the aerodynamic force and moment coefficients and the damping coefficients, are initially obtained using traditional system identification methods via experiments, specifically polynomial regression methods not exceeding the fourth order. These identified uncertainty parameters, most likely not sufficiently accurate, are then used as the initial values for the first phase. In the second phase, the initial values for the network parameters are initialized using the Xavier uniform distribution (Glorot uniform initialization) method.
The 4th-order Runge Kutta and Adam are selected as the ODE solver and the optimizer, respectively. The step size was consistent with the experimental data, set at 1/60 seconds. The length of the prediction horizon $N$ corresponds to the length of the collected experimental data. Through trial and error, we set $N_{1}=N_{2}=10$ for the number of training epochs in each phase of ABNODE. For KNODE and BNODE, 10 epochs were also sufficient, while NODE required 100 epochs to converge. The performance of all models was evaluated using the loss function specified in Eq. (27). We used the weight matrix $\bm{W}$ for normalizing the data and eliminating dimensional units when calculating the loss. Data normalization is performed using Min-Max Scaling.
@@ V-A Accuracy
We compared the prediction loss between ABNODE and other models. Figure 3 illustrates the heat maps of the averaged prediction loss across all the test trials concerning the control input configurations. We observe that ABNODE consistently outperforms the first-principle model, achieving a significant $61.61\%$ reduction in loss across all control input configurations. Furthermore, ABNODE demonstrates improvements of $50.10\%$ and $19.53\%$ in loss reduction compared to SINDYc and BNODE, respectively. NODE performs worse than the first-principle model, with an average loss that is 44.97% higher. Compared to KNODE, BNODE exhibits a 39.21% decrease in the average loss. Therefore, modeling only the dynamics is more effective than modeling both dynamics and kinematics. In subsequent experiments, we primarily compare the performance of ABNODE against the first-principle model, SINDYc, and BNODE.
Inferior model performance is observed with the configuration $(F_{l},F_{r})=(1.7\text{gf},5.4\text{gf})$ due to its longer experimental trajectories, averaging $41.2$ seconds for training and $41.8$ seconds for testing, compared to 31.8 and 34.0 seconds for other configurations. This results in higher cumulative errors from extended prediction horizons and more ODE iterations.
We adjusted the loss and designed two performance metrics, i.e., the time-dependent mean squared error (MSE) and time-dependent cumulative MSE (CMSE) of the prediction loss to delve deeper into the modeling accuracy over specific time periods. Let $\text{MSE}(t_{i})$ denote the model prediction error at time $t_{i}$ and $\text{CMSE}(t_{k})$ denote the sum of MSEs over $k$ steps, i.e.,
@@
where $n$ represents the dimension of system states, and $\bm{x}(t_{i})$ and $\bm{x}_{\text{p}}(t_{i})$ represent the experimental data and model prediction of the states, respectively. The weight matrix $\bm{W}$ in Eq. (30) is the same as in Eq. (27). As an illustration, Fig. 4 showcases the trajectory of CMSE and the distributions of MSE for the configuration $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(0\text{cm},1.2\text{gf},5.4\text{gf})$. ABNODE exhibits the highest modeling accuracy compared to other models, evident from its smallest average MSE, indicating its temporal consistency and robustness.
@@ V-B Generalizability
We evaluated the generalization capability of the four modeling methods using the spiraling experimental dataset. To do so, we trained models using one control input configuration and tested them on neighboring configurations For instance, the configuration $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(1\text{cm},1.2\text{gf},6.1\text{gf})$ had four neighboring configurations, as depicted in Fig. 3, including $(\Delta_{\overline{\bm{r}}_{x}},F_{l},F_{r})=(1\text{cm},1.6\text{gf},5.5\text{gf})$, $(1\text{cm},1.7\text{gf},6.1\text{gf})$, $(0\text{cm},1.2\text{gf},6.1\text{gf})$, and $(2\text{cm},1.2\text{gf},6.1\text{gf})$. We excluded the edge configurations and selected 12 setups, each with four neighboring configurations, resulting in a total of 48 configurations for evaluating the model generalization capability.
Figure 5: Boxplot of the model prediction loss (Eq. 27) across four comparison models. The box represents the interquartile range (IQR), spanning from the first quartile (Q1) to the third quartile (Q3). The line inside the box represents the median while the red diamond dot represents the mean. (Overlap) Beeswarm scatter plot of loss values. Black crosses represent the 48 sets of test results corresponding to each model. The horizontal axis employs a logarithmic scale.